Area of triangle is 40 cms and the length is more than 2 units of its base. What is the length of the triangle?

Answer by checkley79(3341)

L=W+2
AREA=2L+2W
40=2(W+2)+2W
40=2W+4+2W
40=4W+4
5W=40-4
4W=36
W=36/4
W=9 CM. FOR THE WIDTH.
L=9+2=11 CM. FOR THE LENGTH.
PROOF:
40=2*11+2*9
40=22+18
40=40